The width of a rectangular computer screen is 2 inches more than its height.  If the area of the screen is 100 square inches, determine the dimensions of the computer screen
Answer by Boreal(15235)

height is x
width is x+2
A=x(x+2)=100
x^2+2x-100=0
x=(1/2)(-2+/- sqrt (4+400); sqrt (404)=20.1
x=(1/2)(18.1)=9.05 in, only positive root.
x+2=11.05 in
product is 100.00 in^2.
9.05in x 11.05in